What is the equation of the circle having a diameter with endpoints (-4,4) and (2,8)? Please help:)
Answer by scott8148(6628)

the center is midway between the two points at [(-4+2)/2,(4+8)/2] or (-1,6)

the radius is the distance from the center to one of the endpoints
__ r=sqrt[(-1-2)^2+(6-8)^2] __ r=sqrt(9+4) __ r=sqrt(13)

the equation is (x-(-1))^2+(y-6)^2=[sqrt(13)]^2 __ (x+1)^2+(y-6)^2=13